# In this paper, we show that the constant-dimensional Weisfeiler–Leman algorithm for groups (Brachter & Schweitzer, LICS 2020) can be fruitfully used to improve parallel complexity upper bounds on isomorphism testing for several families of groups. In particular, we show:

### v6

@@ 5.2 Additional preliminaries
A group $G$ is the (internal) central product of subgroups $G_{1},G_{2}\leq G$, provided that $G=G_{1}G_{2}$ and $[G_{1},G_{2}]=\{1\}$.
@@
Lemma 5.19 (See, e.g., [Rob82, 3.8.3]).
Let $G=G_{1}\times\ldots\times G_{m}=H_{1}\times\ldots\times H_{n}$ be two direct decompositions of $G$ into directly indecomposable factors. Then $n=m$, and there exists a permutation $\sigma\in\text{Sym}(n)$ such that for all $i$, $G_{i}\cong H_{\sigma(i)}$ and $G_{i}Z(G)\cong H_{\sigma(i)}Z(G)$.
By the preceding lemma, the multiset of subgroups $\{\!\!\{G_{i}Z(G)\}\!\!\}$ is invariant under automorphism.
@@ 5.3 General case
For a group $G$, the non-commuting graph $X_{G}$ has vertex set $G\backslash Z(G)$, [^4] We have changed our definition of the non-commuting graph relative to that in [BS22] to only have vertices corresponding to elements outside the center, in agreement with [AAM06] and prior literature on the non-commuting graph going back to when Erdős first defined it; the rest of the results go through with this modified definition as stated. and an edge $\{g,h\}$ precisely when $[g,h]\neq 1$.
@@
    $G=N_{1}\cdots N_{m}$ is a central decomposition of $G$. For all $i$, $Z(G)\leq N_{i}$ and $N_{i}$ is non-Abelian. In particular, $M$ generates $G$ [^6] We clarify a minor confusion in the definitions and proof from [BS22]. In the proof of their 6.20(2), they write that by definition the process of defining $M$ doesn’t stop until $G=\langle M\rangle$, but they also say that by construction, no central elements are added to $M$, despite their definition allowing central elements in the construction of $M_{i}$ for $i>1$. We believe only allowing non-central elements in the construction of $M_{i}$ to be correct and make the rest of the results go through; in particular, we cannot allow central elements in $M_{i}$ for $i\geq 1$ or some of the “non-abelian components” would actually be cyclic subgroups of the center. Nonetheless, although we do not get $G=\langle M\rangle$ by definition/construction anymore, their argument that $M_{i}=M_{i}Z(G)$ for all $i$ is still correct, and thus $Z(G)\leq\langle M_{1}\rangle\leq\langle M\rangle$, and the process does stop by definition once $G\backslash\langle M_{i}\rangle\leq Z(G)$. Therefore indeed when the process stops we have $G=\langle M\rangle$.
@@
We note that Lem. 5.26 (b)–(c) are purely group-theoretic statements. For our purposes, it is necessary, however, to adapt Lem. 5.26 (a) to control for rounds. This is our second use of our Rank Lemma 4.3, this time applied to the set $M_{i}$ from Definition 5.25.
@@
Let $M_{i,G},M_{G}$ be the sets in $G$ as in Definition 5.25, and let $M_{i,H},M_{H}$ be the corresponding sets in $H$. We show that $f(M_{i,G})=M_{i,H}$ and $f(\langle M_{i,G}\rangle)=\langle M_{i,H}\rangle$ for all $i$. These statements imply that $f(M_{G})=M_{H}$. The proof proceeds by induction over $i$, and within each $i$, we use the Rank Lemma 4.3 applied to $M_{i}$-rank. Note that each $M_{i}$ is closed under taking inverses, since $C_{G}(g)=C_{G}(g^{-1})$.
@@
We first note that if $|C_{G}(g)|\neq|C_{H}(f(g))|$, then Duplicator may win with $2$ pebbles and $2$ rounds. Without loss of generality, suppose that $|C_{G}(g)|>|C_{H}(f(g))|$. Spoiler pebbles $g\mapsto f(g)$. Let $f^{\prime}:G\to H$ be the bijection that Duplicator selects at the next round. Now there exists $x\in C_{G}(g)$ such that $f^{\prime}(x)\not\in C_{H}(f(g))$. Spoiler pebbles $x\mapsto f^{\prime}(x)$ and wins immediately.
Thus $M_{1,G}$ is identified by $(2,2)$-WL. By the Rank Lemma 4.3 applied to $M_{1}$-rank, we get that $f(\langle M_{1,G}\rangle)=\langle M_{1,H}\rangle$ or Spoiler can win with $3$ pebbles in $O(\log n)$ rounds.
As Duplicator must select bijections $f:G\to H$ where $f(\langle M_{1,G}\rangle)=\langle M_{1,H}\rangle$, we may iterate on the above argument replacing $1$ with $i$, to obtain that $f(M_{i,G})=M_{i,H}$ and $f(\langle M_{i,G}\rangle)=\langle M_{i,H}\rangle$. The result now follows by induction. ∎
@@
Let $G$ be non-Abelian, and let $G=G_{1}\times\cdots\times G_{d}$ be an indecompsable direct decomposition. For all $x\in G$, we have a central decomposition $G=C_{x}N_{x}$, with $Z(G)\leq C_{x}\cap N_{x}$. The decomposition is directly induced (Def. 5.20) if and only if $x$ is full for a collection of direct factors of $G$.
@@
Now, by the same argument as above, Duplicator’s bijection $f^{\prime\prime}$ must map $C_{y}$ to $C_{y^{\prime}}$ or Spoiler wins with 1 more pebble (Duplicator can reuse the two pebbles on $x$ and $z$ to implement the strategy of Lem. 5.27) and $O(\log n)$ more rounds. Since $C_{y^{\prime}}\geq C_{f(x)}$, we must have $C_{y}\geq C_{x}$. But then by maximlaity of $C_{x}$, we have $C_{y}=C_{x}$. Therefore $z$ is in $C_{x}$ while $z^{\prime}$ does not commute with $f(x)$, so Spoiler wins immediately. ∎
@@
Let $k\geq 5$, and let $r:=r(n)$. Let $G=G_{1}\times\cdots\times G_{d}$ be a decomposition into indecomposable direct factors. If $G$ and $H$ are not distinguished by $(k,r+O(\log n))$-WL Version II, then there exist indecomposable direct factors $H_{i}\leq H$ such that $H=H_{1}\times\cdots\times H_{d}$; and for all $i\in[d]$, $G_{i}$ and $H_{i}$ are not distinguished by $(k-1,r)$-WL Version II. Furthermore, $G$ and $H$ have isomorphic maximal Abelian direct factors, and $(k-1,r)$-WL Version II fails to distinguish $G_{i}Z(G)$ from $H_{i}Z(H)$, for all $i\in[d]$.
@@
We may assume that $H$ is non-Abelian as well, or Spoiler can win with 2 pebbles in 2 rounds, by pebbling a pair of non-commuting elements of $G$. Let $f:G\to H$ be the bijection that Duplicator selects. By Cor. 5.33, we may assume that $f(\mathcal{F}_{G})=\mathcal{F}_{H}$. It follows that $H$ must admit a decomposition $H=H_{1}\times\ldots\times H_{\ell}$, where the $H_{j}$ factors are directly indecomposable and $\bigcup_{h\in\mathcal{F}_{H}}N_{h}=\bigcup_{j}H_{j}Z(H)\subseteq H$, which we again note is indistinguishable from $\bigcup_{i}G_{i}Z(G)$. Let $X_{G}$ be the non-commmuting graph of $G$, and let $X_{H}$ be the non-commuting graph of $H$. Recall from [AAM06, Proposition 2.1] that as $G,H$ are non-Abelian, $X_{G}$ and $X_{H}$ are connected.
@@
We note that the claims in the preceding paragraph applies to $H$ as well. So if $(k,r+O(\log n))$-WL Version II does not distinguish $G$ and $H$, there must exist a bijection between the connected components of $X_{G}[\bigcup_{i}G_{i}Z(G)]$ and $X_{H}[\bigcup_{i}H_{i}Z(G)]$. Namely, we may assume that $G$ and $H$ admit a decomposition into $\ell=d$ directly indecomposable factors, and that these subgroups are indistinguishable by $(k-1,r)$-WL: for we have a correspondence (after an appropriate reordering of the factors) between $G_{i}Z(H)$ and $H_{i}Z(H)$, where $G_{i}Z(H)$ and $H_{i}Z(H)$ are not distinguished by $(k,r)$-WL. By Lem. 5.15, we have that $(k-1,r)$-WL Version II does not distinguish $G_{i}$ from $H_{i}$. By Thm. 5.12, $G$ and $H$ must have isomorphic maximal Abelian direct factors. So when $G_{i},H_{i}$ are Abelian, we even have $G_{i}\cong H_{i}$. ∎

### v7

@@ 5.2 Additional preliminaries
A group $G$ is the (internal) central product of subgroups $G_{1},G_{2}\leq G$, provided that $G=G_{1}G_{2}$ and $G_{1},G_{2}]=\{1\}$.
@@
Lemma 5.19 (See, e.g., [Rob82, 3.3.8]).
Let $G=G_{1}\times\ldots\times G_{m}=H_{1}\times\ldots\times H_{n}$ be two direct decompositions of $G$ into directly indecomposable factors. Then $n=m$, and there exists a permutation $\sigma\in\text{Sym}(n)$ such that for all $i$, $G_{i}\cong H_{\sigma(i)}$ and $G_{i}Z(G)=H_{\sigma(i)}Z(G)$.
By the preceding lemma, the multiset of subgroups $\{\!\!\{G_{i}Z(G)\}\!\!\}$ is invariant under automorphisms of $G$.
@@ 5.3 General case
[…]
@@
    $G=N_{1}\cdots N_{m}$ is a central decomposition of $G$. For all $i$, $Z(G)\leq N_{i}$ and $N_{i}$ is non-Abelian. In particular, $M$ generates GG. [^6] We clarify a minor confusion in the definitions and proof from [BS22]. In the proof of their 6.20(2), they write that by definition the process of defining $M$ doesn’t stop until $G=\langle M\rangle$, but they also say that by construction, no central elements are added to $M$, despite their definition allowing central elements in the construction of $M_{i}$ for $i>1$. We believe only allowing non-central elements in the construction of $M_{i}$ to be correct and make the rest of the results go through; in particular, we cannot allow central elements in $M_{i}$ for $i\geq 1$ or some of the “non-abelian components” would actually be cyclic subgroups of the center. Nonetheless, although we do not get $G=\langle M\rangle$ by definition/construction anymore, their argument that $M_{i}=M_{i}Z(G)$ for all $i$ is still correct, and thus $Z(G)\leq\langle M_{1}\rangle\leq\langle M\rangle$, and the process does stop by definition once $G\backslash\langle M_{i}\rangle\leq Z(G)$. Therefore indeed when the process stops we have $G=\langle M\rangle$.
@@
We note that Lem. 5.26 (b)–(c) are purely group-theoretic statements. For our purposes, it is necessary, however, to adapt Lem. 5.26 (a) to control for rounds. This is our second use of our Rank Lemma 4.3, this time applied to the set $M_{i}$ from Definition 5.25. We in fact use the Rank Lemma a few times as stated, and in the course of the proof we prove a “relative modulo $\langle M_{i-1}\rangle$” version of the Rank Lemma as well; although such a relative version of the rank lemma modulo a characteristic subgroup would be interesting to state in general, we use several features specific to the case at hand that dissuade us from doing so here.
@@
Let $M_{i,G},M_{G}$ be the sets in $G$ as in Definition 5.25, and let $M_{i,H},M_{H}$ be the corresponding sets in $H$. We show that $f(M_{i,G})=M_{i,H}$ and $f(\langle M_{i,G}\rangle)=\langle M_{i,H}\rangle$ for all $i$. These statements imply that $f(M_{G})=M_{H}$. The proof proceeds by induction over $i$ to show that if the preceding does not hold, then Spoiler can win with 3 pebbles in $\log|\langle M_{i,G}\rangle|)+O(i)$ rounds. Within the inductive step for each $i$, we prove a “quotient” version of the Rank Lemma (Lemma 4.3), applied to $M_{i}$-rank modulo $\langle M_{i-1}\rangle$. Toward that end, we note that each $M_{i}$ is indeed closed under taking inverses, since $C_{G}(g)=C_{G}(g^{-1})$.
@@
We first handle the case of $M_{1}$. If $|C_{G}(g)|\neq|C_{H}(f(g))|$, then Duplicator may win with $2$ pebbles and $2$ rounds. Without loss of generality, suppose that $|C_{G}(g)|>|C_{H}(f(g))|$. Spoiler pebbles $g\mapsto f(g)$. Let $f^{\prime}:G\to H$ be the bijection that Duplicator selects at the next round. Now there exists $x\in C_{G}(g)$ such that $f^{\prime}(x)\not\in C_{H}(f(g))$. Spoiler pebbles $x\mapsto f^{\prime}(x)$ and wins immediately.
Thus $M_{1,G}$ is distinguished by 2 pebbles in 2 rounds. (Since $M_{1}$ is a characteristic subset of any group, and we have shown that for any two finite non-Abelian groups $G,H$, Duplicator must map $M_{1,G}$ to $M_{1,H}$ or Spoiler wins using 2 pebbles in 2 rounds, we may apply the latter statement to the case $G=H$ in order to satisfy the definition of a subset being “distinguished.” We will use this same kind of argument in the future without further comment.) By the Rank Lemma 4.3 applied to $M_{1}$-rank, we get that the subgroup $\langle M_{1,G}\rangle$ is distinguished by 2 pebbles in $\log|\langle M_{1,G}\rangle|+O(1)$ rounds (if Duplicator ever selects a bijection $f$ with $f(\langle M_{1,G}\rangle)\neq\langle M_{1,H}\rangle$, Spoiler can win with 3 pebbles in $\log|\langle M_{1,G}\rangle|+O(1)$ rounds).
Now let $i>1$, and suppose that $M_{i-1,G}$ and $\langle M_{i-1,G}\rangle$ are both distinguished by 2 pebbles in at most $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ rounds for some constant $c$. (Here we give $c$ a name rather than hiding it with big-Oh notation, because we will have various $O(\cdot)$’s floating around and want to keep careful track of which are which.)
First we claim that $M_{i,G}$ is distinguished by 2 pebbles in $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ (n.b. the subscript i1i-1) rounds. Suppose the pebble game is started from the position $g\mapsto h$ for some $g\in M_{i,G}$ and $h\notin M_{i,H}$. If $h\in\langle M_{i-1,H}\rangle$, then by reversing the roles of $G$ and $H$ and using our inductive assumption, we get that Spoiler can win with 2 pebbles in $\log|\langle M_{i-1,H}\rangle|+c(i-1)$ rounds. If $h\notin\langle M_{i-1,H}\rangle$, then we have $h$ is not in $M_{i,H}$ iff $|C_{H}(h)|\neq|C_{G}(g)|$. In the latter case, as above, Spoiler can win with 2 pebbles and 2 rounds. Thus in either case, as long as $c\geq 2$ (which we can indeed choose), Spoiler can win with 2 pebbles in $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ rounds, and therefore $M_{i,G}$ is distinguished using 2 pebbles in $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ rounds.
In the remainder of the proof we will show that the subgroup $\langle M_{i,G}\rangle$ is distinguished by 2 pebbles in $\log|\langle M_{i,G}\rangle|+ci$ rounds.777 We note that if we applied the Rank Lemma here directly, we would instead get 2 pebbles and $\log|\langle M_{i,G}\rangle|+O(1)$ plus the rounds to distinguish $\langle M_{i-1,G}\rangle$, and when applied inductively this would lead to a bound of $O(\log^{2}n)$ instead of $O(\log n)$ rounds. For this part of the proof, we define the relative rank of $g\in G$ as the minimum $r$ such that $g$ can be written as $g=g_{1}g_{2}\dotsc g_{r}m$ where $g_{j}\in M_{i,G}$ (just the set, not $\langle M_{i,G}\rangle$) for all $j=1,\dotsc,r$ and $m\in\langle M_{i-1,G}\rangle$ (the subgroup, not just the set). We define elements of $\langle M_{i-1,G}\rangle$ to have relative rank 0. For $g\notin\langle M_{i,G}\rangle$ we define the relative rank to be $\infty$; it is readily seen that an element has finite relative rank iff it is in $\langle M_{i,G}\rangle$. We use the notation $\text{relrk}(g)$ for the relative rank of $g$. We define relative rank in $H$ analogously, using $M_{i,H}$ and $M_{i-1,H}$ in place of $M_{i,G}$ and $M_{i-1,G}$, respectively; we re-use the same notation $\text{relrk}(h)$ for relative rank in $H$.
We will need the following two facts about relative rank:
  1. R1
    For all $x,y\in G$ (resp. $H$), we have $\text{relrk}(xy)\leq\text{relrk}(x)+\text{relrk}(y)$. If either relative rank is infinite the conclusion is vacuously true. Otherwise, suppose $x=x_{1}\dotsb x_{r}m$ and $y=y_{1}\dotsb y_{s}m^{\prime}$ with $m,m^{\prime}\in\langle M_{i-1}\rangle$; $x_{j},y_{j}\in M_{i}$ for all $j$; and $r=\text{relrk}(x),s=\text{relrk}(y)$. Then $xy=x_{1}\dotsb x_{r}y_{1}\dotsb y_{s}(y^{-1}my)m^{\prime}$, where we have $y^{-1}my$ is in $\langle M_{i-1}\rangle$ since the latter is a normal subgroup.
  2. R2
    For all $x\in\langle M_{i}\rangle$, we have $\text{relrk}(x)\leq|\langle M_{i}\rangle|/|\langle M_{i-1}\rangle|$. (This is where the crucial gains come from, as the $M_{i}$-rank could in principle be as large as $|\langle M_{i}\rangle|$; see also Footnote 7.) For suppose we have $x=x_{1}\dotsb x_{r}m$ with $m\in\langle M_{i-1}\rangle$ and $r>|\langle M_{i}\rangle|/|\langle M_{i-1}\rangle|$. Then among the partial products $x_{1},x_{1}x_{2},x_{1}x_{2}x_{3},\dotsc,x_{1}\dotsb x_{r}$, some coset of $\langle M_{i-1}\rangle$ is encountered twice; suppose that $x_{1}\dotsb x_{j}$ and $x_{1}\dotsb x_{k}$ with $j<k$ are in the same coset of $\langle M_{i-1}\rangle$. Then we can rewrite $x_{1}\dotsb x_{k}$ as $x_{1}\dotsb x_{j}m^{\prime}$ for some $m^{\prime}\in\langle M_{i-1}\rangle$. Finally, since $\langle M_{i-1}\rangle$ is normal, we then get $m^{\prime}x_{k+1}\dotsb x_{r}=x_{k+1}\dotsb x_{r}m^{\prime\prime}$ for a unique $m^{\prime\prime}\in\langle M_{i-1}\rangle$, and thus
    $x=x_{1}\dotsb x_{k}x_{k+1}\dotsb x_{r}m=(x_{1}\dotsb x_{j}m^{\prime})x_{k+1}\dotsb x_{r}m=x_{1}\dotsb x_{j}x_{k+1}\dotsb x_{r}m^{\prime\prime}m,$
    so $\text{relrk}(x)<r$.
Returning to our pebble game starting from the position $g\mapsto h$, without loss of generality (by swapping the roles of $G$ and $H$ if needed), we may assume that $\text{relrk}(g)<\text{relrk}(h)$. If $\text{relrk}(g)=0$ (i. e., $g\in\langle M_{i-1,G}\rangle$) then by the inductive hypothesis, then Spoiler can win with 2 additional pebbles in $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ rounds.
On the other hand, if $\text{relrk}(g)>0$, let $f\colon G\to H$ be the bijection that Duplicator selects at the next round. For brevity let $q=\text{relrk}(g)$. Note that this case works even if $\text{relrk}(h)=\infty$, for then by our assumption that $\text{relrk}(g)<\text{relrk}(h)$, we still have $q=\text{relrk}(g)$ is finite. Now write $g=g_{1}\cdots g_{q}m$, where for each $j$, $g_{j}\in M_{i,G}$ and we have $m\in\langle M_{i-1,G}\rangle$. For $1\leq j\leq k\leq q$, write $g[j,\ldots,k]:=g_{j}\cdots g_{k}$; continuing this notation, write $g[j,\ldots,q+1]=g_{j}g_{j+1}\dotsc g_{q}m$ (that is, we essentially consider $m$ to be at index $q+1$). We consider the following cases.
  • Case 1: Suppose first that for all $x\in G$ with $\text{relrk}(x)<q$, we have $\text{relrk}(x)=\text{relrk}(f(x))$. In this case, Spoiler pebbles $g[2,\ldots,q+1]\mapsto f(g[2,\ldots,q+1])$. Let $f^{\prime}:G\to H$ be the bijection that Duplicator selects at the next round; Spoiler now pebbles $g_{1}\mapsto f^{\prime}(g_{1})$.
    If $\text{relrk}(g_{1})=\text{relrk}(f^{\prime}(g_{1}))=1$, then $f^{\prime}(g_{1})\cdot f(g[2,\ldots,q+1])\neq h$, since $\text{relrk}(f^{\prime}(g_{1}))=1$ and $\text{relrk}(f(g[2,\ldots,q+1]))\leq q-1$, so their product has rank at most $q=\text{relrk}(g)<\text{relrk}(h)$ by fact R1 above. Spoiler now wins since the map on the pebbled elements $(g,g[2,\ldots,q+1],g_{1})\mapsto(h,f(g[2,\ldots,q+1]),f^{\prime}(g_{1}))$ does not satisfy the winning condition for Spoiler in the Version I (and hence, the Version II) game.
    On the other hand, suppose instead that $1=\text{relrk}(g_{1})<\text{relrk}(f^{\prime}(g_{1}))$. Note that $\text{relrk}(f^{\prime}(g_{1}))>1$ implies that $f^{\prime}(g_{1})$ is not in $M_{i,H}$ (relative rank is $\leq 1$ iff an element is in $M_{i,H}\cdot\langle M_{i-1,H}\rangle$). But since $g_{1}$ is, by definition, an element of $M_{i,G}$, from the fact that $M_{i,G}$ is distinguished, we have that Spoiler can win with $2$ additional pebbles and $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ additional rounds. Note that only 2 pebbles are needed rather than 3, since the original pebble on $g$ can be reused at this point.
  • Case 2: Suppose instead that the hypothesis of Case 1 is not satisfied. Then there exists some $x\in\langle M_{i,G}\rangle$ with $\text{relrk}(x)\neq\text{relrk}(f(x))$ and $\text{relrk}(x)<q=\text{relrk}(g)$. We will now do a binary-search-like procedure to eventually reach either (a) an element of $M_{i,G}$ that is mapped outside of $M_{i,H}$ or (b) an element of $\langle M_{i-1,G}\rangle$ that is mapped outside of $\langle M_{i-1,H}\rangle$. To make this clearer, we introduce two additional parameters as in binary search. Let $lo=1$ and $hi=q+1$. In the next two rounds, Spoiler pebbles $x:=g[lo,\ldots,lo+\lceil(hi-lo)/2\rceil-1]$ and $y:=g[lo+\lceil(hi-lo)/2\rceil,\ldots,hi]$. Note that $\text{relrk}(x)=\lceil(hi-lo)/2\rceil$ and $\text{relrk}(g)=\text{relrk}(x)+\text{relrk}(y)$, for if $x$ or $y$ could be written as a shorter word in $M_{i,G}$, then so could $g$, by fact R1.
    Let $f^{\prime}:G\to H$ be the next bijection that Duplicator selects. If $h\neq f^{\prime}(x)f^{\prime}(y)$ then Spoiler immediately wins. On the other hand, if $h=f^{\prime}(x)f^{\prime}(y)$, then either
    $\text{relrk}(x)<\text{relrk}(f^{\prime}(x))\qquad\text{ or }\qquad\text{relrk}(y)<\text{relrk}(f^{\prime}(y)),$
    since $\text{relrk}(h)>\text{relrk}(g)=\text{relrk}(x)+\text{relrk}(y)$.
    Without loss of generality, suppose that $\text{relrk}(x)<\text{relrk}(f^{\prime}(x))$. Spoiler then updates $hi$ to $lo+\lceil(hi-lo)/2\rceil-1$, and as $x\mapsto f^{\prime}(x)$ is already pebbled, Spoiler iterates on this strategy as in binary search, reusing the pebbles on $g$ and $y$ in the next iteration, with $x$ now playing the role that $g$ played in the first iteration. Thus, iterating this procedure uses at most $3$ pebbles (the one that started on $g$ plus 2 more). After at most $\log q+O(1)$ iterations, we reach the case where $hi=lo+1$, and we have pebbled either (a) an element $z\mapsto f^{\prime\prime}(z)$ such that $z\in M_{i,G}$ and $\text{relrk}(f^{\prime\prime}(z))>1$ or (b) an element $m\mapsto f^{\prime\prime}(m)$ with $m\in\langle M_{i-1,G}\rangle$ and $\text{relrk}(f^{\prime\prime}(m))>0$. At this point, there are two pebbles in play that can be reused.
    By fact R2, every element of $\langle M_{i,G}\rangle$ has relative rank at most $|\langle M_{i,G}\rangle|/|\langle M_{i-1,G}\rangle|$, and thus only $\log\frac{|\langle M_{i,G}\rangle|}{|\langle M_{i-1,G}\rangle|}+O(1)$ rounds of the preceding iteration are required. Choose $c$ large enough so that it is greater than the preceding $O(1)$.
    Finally, since both $M_{i,G}$ and $\langle M_{i-1,G}\rangle$ are distinguished by 2 pebbles in $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ rounds, and we have either (a) pebbled an element of $M_{i,G}$ mapped outside of $M_{i,H}$ or (b) pebbled an element of $\langle M_{i-1,G}\rangle$ mapped outside of $\langle M_{i-1,H}\rangle$, in either case Spoiler can reuse the aforementioned 2 pebbles to win in $\log|\langle M_{i-1,G}\rangle|+c(i-1)$ additional rounds.
    Thus $\langle M_{i,G}\rangle$ is distinguished by 2 pebbles in at most
    $\log\frac{|\langle M_{i,G}\rangle|}{|\langle M_{i-1,G}\rangle|}+c+\log|\langle M_{i-1,G}\rangle|+c(i-1)$
    total rounds. The preceding expression readily simplifies to $\log|\langle M_{i,G}\rangle|+ci$, as claimed.
Note that, since the $M_{i}$ are constructed so that $\langle M_{i}\rangle$ strictly contains $\langle M_{i-1}\rangle$, the maximum value of $i$ is at most $\log|G|=\log n$. By the above induction, we then get that each $M_{i}$ is distinguished by 2 pebbles $\log|\langle M_{i}\rangle|+ci$ rounds, where the latter is at most $\log n+c\log n=O(\log n)$. Finally, we apply the Rank Lemma 4.3 to get that $M_{i}$-rank must be preserved or Spoiler can win with 3 pebbles and $O(\log n)+\log n+O(1)=O(\log n)$ rounds. ∎
@@
Let $G$ be non-Abelian, and let $G=G_{1}\times\cdots\times G_{d}$ be an indecomposable direct decomposition. For all $x\in G$, we have a central decomposition $G=C_{x}N_{x}$, with $Z(G)\leq C_{x}\cap N_{x}$. The decomposition is directly induced (Def. 5.20) if and only if $x$ is full for a collection of direct factors of $G$.
@@
Now, by the same argument as above, Duplicator’s bijection $f^{\prime\prime}$ must map $C_{y}$ to $C_{y^{\prime}}$ or Spoiler wins with 1 more pebble (Duplicator can reuse the two pebbles on $x$ and $z$ to implement the strategy of Lem. 5.27) and $O(\log n)$ more rounds. Since $C_{y^{\prime}}\geq C_{f(x)}$, we must have $C_{y}\geq C_{x}$. But then by maximality of $C_{x}$, we have $C_{y}=C_{x}$. Therefore $z$ is in $C_{x}$ while $z^{\prime}$ does not commute with $f(x)$, so Spoiler wins immediately. ∎
@@
Let $k\geq 5$, and let $r:=r(n)$. Let $G=G_{1}\times\cdots\times G_{d}$ be a decomposition into indecomposable direct factors. If $G$ and $H$ are not distinguished by $(k,r+O(\log n))$-WL Version II, then there exist indecomposable direct factors $H_{i}\leq H$ such that $H=H_{1}\times\cdots\times H_{d}$; and for all $i\in[d]$, $G_{i}$ and $H_{i}$ are not distinguished by (k1,r)(k-1,r)-WL Version II. Furthermore, $G$ and $H$ have isomorphic maximal Abelian direct factors, and (k1,r)(k-1,r)-WL Version II fails to distinguish $G_{i}Z(G)$ from $H_{i}Z(H)$, for all $i\in[d]$.
@@
We may assume that $H$ is non-Abelian as well, or Spoiler can win with 2 pebbles in 2 rounds, by pebbling a pair of non-commuting elements of $G$. Let $f:G\to H$ be the bijection that Duplicator selects. By Cor. 5.33, we may assume that $f(\mathcal{F}_{G})=\mathcal{F}_{H}$. It follows that $H$ must admit a decomposition $H=H_{1}\times\ldots\times H_{\ell}$, where the $H_{j}$ factors are directly indecomposable and $\bigcup_{h\in\mathcal{F}_{H}}N_{h}=\bigcup_{j}H_{j}Z(H)\subseteq H$, which we again note is indistinguishable from $\bigcup_{i}G_{i}Z(G)$. Let $X_{G}$ be the non-commuting graph of $G$, and let $X_{H}$ be the non-commuting graph of $H$. Recall from [AAM06, Proposition 2.1] that as $G,H$ are non-Abelian, $X_{G}$ and $X_{H}$ are connected.
@@
We note that the claims in the preceding paragraph applies to $H$ as well. So if $(k,r+O(\log n))$-WL Version II does not distinguish $G$ and $H$, there must exist a bijection between the connected components of $X_{G}[\bigcup_{i}G_{i}Z(G)]$ and $X_{H}[\bigcup_{i}H_{i}Z(G)]$. Namely, we may assume that $G$ and $H$ admit a decomposition into $\ell=d$ directly indecomposable factors, and that these subgroups are indistinguishable by (k1,r)(k-1,r)-WL: for we have a correspondence (after an appropriate reordering of the factors) between $G_{i}Z(H)$ and $H_{i}Z(H)$, where $G_{i}Z(H)$ and $H_{i}Z(H)$ are not distinguished by $(k,r)$-WL. By Lem. 5.15, we have that (k1,r)(k-1,r)-WL Version II does not distinguish $G_{i}$ from $H_{i}$. By Thm. 5.12, $G$ and $H$ must have isomorphic maximal Abelian direct factors. So when $G_{i},H_{i}$ are Abelian, we even have $G_{i}\cong H_{i}$. ∎